OpenStudy (anonymous):

Write a polynomial of least degree with roots 4 and -4 - 4.

OpenStudy (anonymous):

4 and -4 are roots/solutions, so you have to do the inverse, which is to write a polynomial. x=4 is (x-4) and x=-4 is (x+4) Multilply (use FOIL): (x+4)(x-4), and that's the polynomial.

OpenStudy (samirmath01):

polynomial is for example: \[ (x-4)(x+4) = x^2+4x-4x-16 = x^2-16\]